OpenStudy (anonymous):

What is the midpoint between (-8, 5) and (2, -2)? (-5, 1.5) (-5, 3.5) (-6, 3) (-3, 1.5)

OpenStudy (saifoo.khan):

Use the midpoint formula

OpenStudy (saifoo.khan):

Here: http://geometry.jdeer.com/Images/c1/midpoint%20formula.gif

OpenStudy (lgbasallote):

\[\Large \frac{x_1 + x_2}{2}, \; \frac{y_1 +y_2}{2}\] \[\Large \frac{-8+2}{2}, \; \frac{5 + -2}{2}\] does that help?

OpenStudy (saifoo.khan):

No, it doesn't. @lgbasallote

OpenStudy (lgbasallote):

yeah...you can do better. im sure

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

-3,1.5 ?